# SKADA-Bench: Benchmarking Unsupervised Domain Adaptation Methods with Realistic Validation

### v1

@@ 3.2 Compared DA methods
In this section, we present the different families of domain adaptation methods that we compare in our benchmark. We group the methods into four categories: reweighting methods, mapping methods, subspace methods, and others. We provide a brief description of each method and the corresponding references.
Reweighting methods.
@@
These methods aim to reweight the source data to make it closer to the target data. The weights are estimated using different methods such as kernel density estimation (Dens. RW) [50], Gaussian estimation (Gauss. RW) [46], discriminative estimation (Discr. RW) [46], or nearest-neighbors (NN RW) [30]. Other reweighting estimate weights by minimizing a divergence between the source and target distributions such as Kullback-Leibler Importance Estimation Procedure (KLIEP) [51] or Kernel Mean Matching (KMM) [21]. Finally, we also include the MMDTarS method [60] that uses a Maximum Mean Discrepancy (MMD) to estimate the weights under the target shift hypothesis.
@@
Mapping methods. These methods aim to find a mapping between the source and target data that minimizes the distribution shift. The Correlation Alignment method (CORAL) [52] aligns the second-order statistics of source and target distributions. The Maximum Mean Discrepancy (MMD-LS) method [60] minimizes the MMD to estimate an affine Location-Scale mapping. Finally, the Optimal Transport (OT) mapping methods [7] use the optimal transport plan to align with a non-linear mapping of the source and target distributions with exact OT (MapOT), entropic regularization (EntOT), or class-based regularization (ClassRegOT). Finally, the Linear OT method [13] uses a linear mapping to align the source and target distributions, assuming Gaussian distributions.
@@
Subspace methods. These methods aim to learn a subspace where the source and target data have the same distribution. The Transfer Component Analysis (TCA) method [39] searches for a kernel embedding that minimizes the MMD divergence between the domains while preserving the variance. The Subspace Alignment (SA) method [12] aims to learn a subspace where the source and target have their covariance matrices aligned. The Transfer Subspace Learning (TSL) method [47] aims to learn a subspace using classical supervised loss functions on the source (e.g., PCA, Fisher LDA) but regularized so that the source and target data have the same distribution once projected on the subspace. Finally, the Joint Principal Component Analysis (JPCA) method is a simple baseline that concatenates source and target data before applying a PCA on all data.
@@
Others. We also include other methods that do not fit into the previous categories. The Domain Adaptation SVM (DASVM) method [3] is a self-labeling method that iteratively updates SVM estimators by adding new target samples with predicted labels and removing source samples. The Joint Distribution Optimal Transport (JDOT) method [6] aims to learn a target predictor that minimizes an OT loss between the joint source and target distributions. The Optimal Transport Label Propagation (OTLabelProp) method [48] uses the optimal transport plan to propagate labels from the source to the target domain.
@@ 3.3 Compared datasets
Simulated datasets.
The objective of the simulated datasets is to evaluate the performance of the DA methods under different types of shifts. Knowing that multiple DA methods have been built to handle specific shifts, evaluating them with this dataset will demonstrate whether they perform as expected and if they are properly validated.
The 4 simulated shifts in 2D, covariate (Cov. shift), target (Tar. shift) conditional (Cond. shift) and Subspace (Sub. shift) shift are illustrated in Figure 1. The source domain is represented by two non-linearly separable classes generated from one large and several smaller Gaussian blobs. In the experiments, the level of noise has been adjusted from Figure 1 to make the problem more difficult. For the subspace shift scenario, the source domain consists of one class represented by a large Gaussian blob and another class comprising Gaussian blobs positioned along the sides of the large one. The target domain is flipped along the diagonal, making the task challenging in the original space but feasible upon diagonal projection.
@@
Office 31 [26] CV Decaff [10] + PCA 6 31 470 $\pm$ 350 100
Office Home [55] CV ResNet [19] + PCA 12 65 3897 $\pm$ 850 100
MNIST/USPS [28] CV Vect + PCA 2 10 3000 / 10000 50
20 Newsgroup [27] NLP LLM [43, 57] + PCA 6 2 3728 $\pm$ 174 50
Amazon Review [32, 33] NLP LLM [43, 57] + PCA 12 4 2000 50
Mushrooms [8] Tabular One Hot Encoding 2 2 4062 $\pm$ 546 117
Phishing [34] Tabular NA 2 2 5527 $\pm$ 1734 30
BCI [54] Biosignals Cov+TS [1] 9 4 288 253
Real-word datasets.
The real-world datasets used in our benchmark are summarized in Table 1. We select $8$ datasets from different modalities and tasks: Computer Vision (CV) with Office31 [26], Office Home [55], and MNIST/USPS [28], Natural Language Processing (NLP) with 20Newsgroup [27] and Amazon Review [33], Tabular Data with Mushrooms [8] and Phishing [34], and Biosignals with BCI Competition IV [54]. The datasets are chosen to represent a wide range of shifts and to evaluate the performance of the methods on different types of data.
The datasets are preprocessed with feature extraction to ensure reasonable performance when trained on each domain. For example, images are embedded using pre-trained models followed by a PCA (except MNIST/USPS where only PCA is used), and textual data is embedded using Large Language Models (LLM) [43, 57] before applying a PCA. The tabular data are one-hot encoded to transform categorical data into numerical data. The biosignals from Brain-Computer Interface (BCI) data are embedded using the state-of-the-art tangent space representation proposed in [1]. The datasets are split into pairs of source and target domains totaling $51$ adaptation tasks in the benchmark. More details about the datasets and pre-processing are available in Appendix B.
@@ 4 Benchmark results
We now present the results of the benchmark. Training and evaluation across all experiments required 1,215 CPU-hours on a standard Slurm [58] cluster. We first discuss and compare the performances of the methods on the different datasets. Then, a detailed study of the unsupervised scorers is provided.
@@ 4.1 Performance of the DA methods
First, we report the realistic performances of the different methods when using their selected scorer on the different datasets in Table 2. The cells showcasing a significant change in performance with the Wilcoxon test are highlighted with colors. Blue indicates an increase in performance, while red indicates a loss. The intensity of the color corresponds to the magnitude of the gain or loss - the darker the shade, the larger the positive or negative change. Cells with a NA values indicate that the method was not applicable to the dataset (DASVM is limited to binary classification) or that the method has reached a timeout. We also report the best scorer and the average rank of the methods for all datasets. In addition to Table 2 providing realistic performance estimations, we also report in Table D.12 (Appendix D) the results when using the non-realistic supervised scorer.
Table 2: Accuracy score for all datasets compared for all the methods for simulated and real-life datasets. The color indicates the amount of the improvement. A white color means the method is not statistically different from Train Src (Train on source). Blue indicates that the score improved with the DA methods, while red indicates a decrease. The darker the color, the more significant the change.
@@
Train Src 0.88 0.85 0.66 0.19 0.59 0.56 0.54 0.59 0.7 0.72 0.91 0.55 9.75
Train Tgt 0.92 0.93 0.82 0.98 0.88 0.8 0.96 1.0 0.73 1.0 0.97 0.64 1.06
Reweighting Dens. RW [50] 0.88 0.86 0.66 0.18 0.57 0.55 0.54 0.58 0.7 0.71 0.91 0.55 IW 10.76
Disc. RW [46] 0.85 0.83 0.71 0.18 0.58 0.53 0.5 0.6 0.68 0.75 0.91 0.56 CircV 11.12
Gauss. RW [46] 0.89 0.86 0.65 0.21 0.2 0.44 0.11 0.54 0.6 0.51 0.46 0.25 CircV 19.42
KLIEP [51] 0.88 0.86 0.66 0.19 0.59 0.56 0.54 0.6 0.69 0.72 0.91 0.55 CircV 10.36
KMM [21] 0.89 0.87 0.64 0.15 0.58 0.55 0.52 0.7 0.57 0.74 0.91 0.52 CircV 12.11
NN RW [30] 0.89 0.86 0.67 0.15 0.58 0.55 0.54 0.59 0.66 0.71 0.91 0.54 CircV 11.91
MMDTarS [60] 0.88 0.86 0.64 0.2 0.56 0.55 0.54 0.59 0.7 0.74 0.91 0.55 IW 9.51
Mapping CORAL [52] 0.66 0.84 0.66 0.19 0.59 0.57 0.62 0.73 0.69 0.72 0.92 0.62 CircV 7.10
MapOT [7] 0.72 0.57 0.82 0.02 0.55 0.51 0.61 0.76 0.67 0.63 0.84 0.47 PE 10.98
EntOT [7] 0.71 0.6 0.82 0.12 0.58 0.58 0.6 0.83 0.62 0.75 0.86 0.54 CircV 9.75
ClassRegOT [7] 0.74 0.58 0.81 0.11 0.61 0.53 0.62 0.96 0.68 0.82 0.88 0.52 IW 8.71
LinOT [13] 0.73 0.73 0.76 0.18 0.59 0.57 0.64 0.82 0.7 0.76 0.91 0.61 CircV 5.33
MMD-LS [60] 0.65 0.68 0.81 0.52 0.55 0.54 0.52 0.97 0.68 0.86 0.88 0.56 IW 9.66
Subspace JPCA 0.88 0.85 0.66 0.15 0.55 0.47 0.51 0.77 0.69 0.78 0.9 0.54 PE 8.77
SA [12] 0.74 0.68 0.8 0.11 0.59 0.57 0.56 0.88 0.66 0.88 0.89 0.53 CircV 8.53
TCA [39] 0.46 0.48 0.55 0.56 0.04 NA 0.11 0.57 0.6 0.45 NA 0.27 CircV 19.57
TSL [47] 0.88 0.85 0.66 0.19 0.59 0.2 0.25 0.68 0.7 0.56 0.86 0.25 IW 14.65
Other JDOT [6] 0.72 0.57 0.82 0.14 0.6 0.51 0.63 0.77 0.67 0.63 0.8 0.46 DEV 10.12
OTLabelProp [48] 0.72 0.59 0.81 0.05 0.61 0.56 0.62 0.86 0.67 0.64 0.86 0.5 CircV 10.49
DASVM [3] 0.89 0.86 0.65 0.14 NA NA NA 0.68 NA 0.78 0.88 NA CircV 11.00
Simulated data with known shifts.
On simulated data with known shifts (i.e., underlined datasets), DA methods tend to show a small but significant gain on the shift they were designed for. However, they rarely reach the Train Tgt performances, and as expected, the methods perform poorly on the shift they were not designed for. For Cov. shift, the improvement with reweighting methods is very limited. We believe that using a complex base estimator (SVM with an RBF kernel) enables us to train an estimator that works well on both source and target, leading to less impressive results. The Appendix D contains detailed tables for two other base estimators (i.e., Logistic Regression in Table D.9 and XGBoost in Table D.11). These tables show that the gain is larger for Logistic Regression, a linear classifier, although the overall performance is lower. Note that in the literature, the base estimator used is often a simple method, which can explain the performance gains reported in various papers.
Real life datasets.
On the real-life datasets, we observe that DA methods do not consistently outperform the Train Src approach. The only exceptions are the LinOT and CORAL methods, which either result in a significant improvement or perform comparably to the Train Src approach. OT-based mapping methods show gains on MNIST/USPS and 20NewsGroups. Other methods, such as ClassRegOT, MMD-LS, JPCA, and SA, occasionally show better accuracy than Train Src. However, interestingly, DA methods fail to improve performance on datasets like Office31, AmazonReview, and Phishing. For the AmazonReview dataset, Train Src performs exceptionally well, possibly because the selected LLM has been partially trained on similar Amazon review data, leading to invariant representations [57].
Looking at the average rank of the methods, we observe that the Train Src approach has a relatively small rank of 9.75 out of 20, while the best-rank DA methods are LinOT with a rank of 5.33 and CORAL with 7.10. Other methods with a lower rank than Train Src are ClassRegOT, MMDTarS, JPCA, and SA.
Selected scorer per DA method.
@@
We observe that the best scorer differs across methods, Circular Validation has been selected 12 out of 20 times as the best scorer, followed by Importance Weighting 5 out of 20 times. Table D.12 in the supplementary material provides the accuracy results with the supervised scorer. It is worth noting that the supervised scorer generally outperforms the unsupervised ones, and several methods significantly outperform Train Src in each dataset. It is crucial to choose the model realistically to avoid producing overly optimistic results, as many data analysis papers have done (see able A.1).

### v2

@@ 3.2 Compared DA methods
In this section, we present the different families of domain adaptation methods that we compare in our benchmark. The shallow methods are grouped into four categories: reweighting methods, mapping methods, subspace methods, and others. For Deep DA methods, we consider three domain invariant feature methods. We provide a brief description of each method and the corresponding references.
@@
Reweighting methods. These methods aim to reweight the source data to make it closer to the target data. The weights are estimated using different methods such as kernel density estimation (Dens. RW) (Sugiyama & Müller, 2005), Gaussian estimation (Gauss. RW) (Shimodaira, 2000), discriminative estimation (Discr. RW) (Shimodaira, 2000), or nearest-neighbors (NN RW) (Loog, 2012). Other reweighting estimate weights by minimizing a divergence between the source and target distributions such as Kullback-Leibler Importance Estimation Procedure (KLIEP) (Sugiyama et al., 2007b) or Kernel Mean Matching (KMM) (Huang et al., 2006). Finally, we also include the MMDTarS method (Zhang et al., 2013) that uses a Maximum Mean Discrepancy (MMD) to estimate the weights under the target shift hypothesis.
@@
Mapping methods. These methods aim to find a mapping between the source and target data that minimizes the distribution shift. The Correlation Alignment method (CORAL) (Sun et al., 2017) aligns the second-order statistics of source and target distributions. The Maximum Mean Discrepancy (MMD-LS) method (Zhang et al., 2013) minimizes the MMD to estimate an affine Location-Scale mapping. Finally, the Optimal Transport (OT) mapping methods (Courty et al., 2017b) use the optimal transport plan to align with a non-linear mapping of the source and target distributions with exact OT (MapOT), entropic regularization (EntOT), or class-based regularization (ClassRegOT). Finally, the Linear OT method (Flamary et al., 2020) uses a linear mapping to align the source and target distributions, assuming Gaussian distributions.
@@
Subspace methods. These methods aim to learn a subspace where the source and target data have the same distribution. The Transfer Component Analysis (TCA) method (Pan et al., 2011) searches for a kernel embedding that minimizes the MMD divergence between the domains while preserving the variance. The Subspace Alignment (SA) method (Fernando et al., 2013) aims to learn a subspace where the source and target have their covariance matrices aligned. The Transfer Subspace Learning (TSL) method (Si et al., 2010) aims to learn a subspace using classical supervised loss functions on the source (e.g., PCA, Fisher LDA) but regularized so that the source and target data have the same distribution once projected on the subspace. Finally, the Joint Principal Component Analysis (JPCA) method is a simple baseline that concatenates source and target data before applying a PCA.
@@
Others. We also include other methods that do not fit into the previous categories. The Domain Adaptation SVM (DASVM) method (Bruzzone & Marconcini, 2010a) is a self-labeling method that iteratively updates SVM estimators by adding new target samples with predicted labels and removing source samples. The Joint Distribution Optimal Transport (JDOT) method (Courty et al., 2017a) aims to learn a target predictor that minimizes an OT loss between the joint source and target distributions. The Optimal Transport Label Propagation (OTLabelProp) method (Solomon et al., 2014) uses the optimal transport plan to propagate labels from the source to the target domain.
Deep DA methods. These methods aim to reduce the divergence between the source and target data distributions within the learned feature space while simultaneously learning a classifier on source data. The training loss consists in a traditional supervised loss on labeled source data and a second term measuring the discrepancy between the source and target distributions. The methods implemented in the Deep DA Benchmark use different discrepancies, such as covariance distance (Sun & Saenko, 2016) for DeepCORAL, adversarial loss (Ganin et al., 2016a) for DANN and optimal transport distance (Damodaran et al., 2018a) for DeepJDOT. Note that these approaches are not part of the main shallow DA benchmark but have been added to provide an interesting comparison of DA performances between shallow and Deep methods on computer vision and biomedical data.
@@ 3.3 Compared datasets
Simulated datasets. The objective of the simulated datasets is to evaluate the performance of the DA methods under different types of shifts. Knowing that multiple DA methods have been built to handle specific shifts, evaluating them with this dataset will demonstrate whether they perform as expected and if they are properly validated.
The four simulated shifts in 2D, covariate (Cov. shift), target (Tar. shift) conditional (Cond. shift) and Subspace (Sub. shift) shift are illustrated in Figure 1. The source domain is represented by two non-linearly separable classes generated from one large and several smaller Gaussian blobs. In the experiments, the level of noise has been adjusted from Figure 1 to make the problem more difficult. For the subspace shift scenario, the source domain consists of one class represented by a large Gaussian blob and another class comprising Gaussian blobs positioned along the sides of the large one. The target domain is flipped along the diagonal, making the task challenging in the original space but feasible upon diagonal projection.
@@
Office 31 (Koniusz et al., 2017) CV Decaff + PCA (Donahue et al., 2014) 6 31 470 $\pm$ 350 100
Office Home (Venkateswara et al., 2017) CV ResNet + PCA (He et al., 2016) 12 65 3897 $\pm$ 850 100
MNIST/USPS (Liao & Carneiro, 2015) CV Vect + PCA 2 10 3000 / 10000 50
20 Newsgroup (Lang, 1995) NLP LLM + PCA (Reimers & Gurevych (2019), Xiao et al. (2023a)) 6 2 3728 $\pm$ 174 50
Amazon Review (McAuley & Leskovec (2013), McAuley et al. (2015)) NLP LLM + PCA (Reimers & Gurevych (2019), Xiao et al. (2023a)) 12 4 2000 50
Mushrooms (Dai et al., 2007) Tabular One Hot Encoding 2 2 4062 $\pm$ 546 117
Phishing (Mohammad et al., 2012) Tabular NA 2 2 5527 $\pm$ 1734 30
BCI (Tangermann et al., 2012) Biosignals Cov+TS (Barachant et al., 2012) 9 4 288 253
Real-word datasets. The real-world datasets used in our benchmark are summarized in Table 1. We select $8$ datasets from different modalities and tasks: Computer Vision (CV) with Office31 (Koniusz et al., 2017), Office Home (Venkateswara et al., 2017), and MNIST/USPS (Liao & Carneiro, 2015), Natural Language Processing (NLP) with 20Newsgroup (Lang, 1995) and Amazon Review (McAuley et al., 2015), Tabular Data with Mushrooms (Dai et al., 2007) and Phishing (Mohammad et al., 2012), and Biosignals with BCI Competition IV (Tangermann et al., 2012). The datasets are chosen to represent a wide range of shifts and to evaluate the performance of the methods on different types of data.
Before using shallow DA methods, the datasets are preprocessed with feature extraction to ensure reasonable performance when trained on each domain. For example, images are embedded using Deep pre-trained models followed by a PCA (except MNIST/USPS where only PCA is used), and textual data is embedded using Large Language Models (LLM) (Reimers & Gurevych, 2019; Xiao et al., 2023a) before applying a PCA. The tabular data are one-hot encoded to transform categorical data into numerical data. The biosignals from Brain-Computer Interface (BCI) data are embedded using the state-of-the-art tangent space representation proposed in Barachant et al. (2012). For Deep DA methods, only 4 datasets are used: Office31, Office Home, MNIST/USPS and BCI. Since these methods focus on learning feature representations, the data are used in their raw form. The datasets are split into pairs of source and target domains totaling $51$ adaptation tasks in the benchmark. More details about the datasets and pre-processing are available in Appendix C.
@@ 4 Benchmark results
We now present the results of the benchmark. Training and evaluation across all shallow experiments required 1,215 CPU-hours on a standard Slurm (Yoo et al., 2003) cluster, while the Deep DA experiments required 244 GPU-hours. We first discuss and compare the performances of the methods on the different datasets. Then, a detailed study of the unsupervised scorers is provided.
@@ 4.1 Performance of the DA methods
First, we report the realistic performances of the different methods when using their selected scorer on the different datasets in Table 2. The cells showcasing a significant change in performance with the Wilcoxon test are highlighted with colors. Blue indicates an increase in performance, while red indicates a loss. The intensity of the color corresponds to the magnitude of the gain or loss - the darker the shade, the larger the positive or negative change. Cells with a NA values indicate that the method was not applicable to the dataset (DASVM is limited to binary classification) or that the method has reached a timeout. We also report the best scorer and the average rank of the methods for all real datasets. In addition to Table 2 providing realistic performance estimations with the best realistic scorer, we also report in Table 20 (Appendix E) the results when using the non-realistic supervised scorer.
Table 2: Accuracy score for all datasets compared for all the shallow methods for simulated and real-life datasets. The color indicates the amount of the improvement. A white color means the method is not statistically different from Train Src (Train on source). Blue indicates that the score improved with the DA methods, while red indicates a decrease. The darker the color, the more significant the change.
@@
Train Src 0.88 0.85 0.66 0.19 0.65 0.56 0.54 0.59 0.7 0.72 0.91 0.55 10.66
Train Tgt 0.92 0.93 0.82 0.98 0.89 0.8 0.96 1.0 0.73 1.0 0.97 0.64 1.55
Reweighting Dens. RW 0.88 0.86 0.66 0.18 0.62 0.56 0.54 0.58 0.7 0.71 0.91 0.55 IW 12.20
Disc. RW 0.85 0.83 0.71 0.18 0.63 0.54 0.5 0.6 0.68 0.75 0.91 0.56 CircV 8.75
Gauss. RW 0.89 0.86 0.65 0.21 0.22 0.44 0.11 0.54 0.55 0.51 0.46 0.25 CircV 16.45
KLIEP 0.88 0.86 0.66 0.19 0.65 0.56 0.54 0.6 0.69 0.72 0.91 0.55 CircV 10.56
KMM 0.89 0.85 0.64 0.16 0.64 0.54 0.52 0.7 0.57 0.74 0.91 0.52 CircV 11.74
NN RW 0.89 0.86 0.67 0.15 0.65 0.55 0.54 0.59 0.66 0.71 0.91 0.54 CircV 9.15
MMDTarS 0.88 0.86 0.64 0.2 0.6 0.56 0.54 0.59 0.7 0.74 0.91 0.55 IW 10.81
Mapping CORAL 0.74 0.7 0.76 0.18 0.65 0.57 0.62 0.73 0.7 0.72 0.92 0.62 CircV 5.08
MapOT 0.72 0.57 0.82 0.02 0.6 0.51 0.61 0.76 0.68 0.63 0.84 0.47 PE 10.21
EntOT 0.71 0.6 0.82 0.12 0.64 0.58 0.6 0.83 0.62 0.75 0.86 0.54 CircV 9.40
ClassRegOT 0.74 0.58 0.81 0.11 NA 0.53 0.62 0.97 0.68 0.82 0.89 0.52 IW 8.25
LinOT 0.73 0.73 0.76 0.18 0.66 0.57 0.64 0.82 0.7 0.76 0.91 0.61 CircV 4.06
MMD-LS 0.78 0.72 0.76 0.56 0.65 0.56 0.55 0.97 0.63 0.85 NA 0.5 MixVal 8.22
Subspace JPCA 0.88 0.85 0.66 0.15 0.62 0.48 0.51 0.77 0.69 0.78 0.9 0.54 PE 8.98
SA 0.74 0.68 0.8 0.11 0.65 0.57 0.56 0.88 0.67 0.78 0.89 0.53 CircV 7.80
TCA 0.52 0.47 0.51 0.62 0.04 0.02 0.07 0.61 0.61 0.49 0.48 0.26 DEV 17.58
TSL 0.88 0.85 0.66 0.2 0.63 0.48 0.45 0.63 0.69 0.45 0.89 0.26 IW 15.09
Other JDOT 0.72 0.58 0.82 0.13 0.6 0.42 0.59 0.79 0.67 0.65 0.79 0.47 IW 11.42
OTLabelProp 0.72 0.59 0.8 0.07 0.66 0.56 0.62 0.86 0.67 0.64 0.86 0.5 CircV 10.01
DASVM 0.89 0.86 0.65 0.15 NA NA NA 0.87 NA 0.83 0.85 NA MixVal 7.29
Simulated data with known shifts. DA methods tend to show a significant gain on the shift they were designed for. It is especially true for mapping methods which greatly outperforms the Train Src approach under conditional shift (Cond. shift), almost reaching the Train Tgt performance for EntOT and ClassRegOT. The results also highlight that the mapping methods struggle with target shift (Tar. shift), which is a well-known limitation of this kind of approach (Redko et al., 2019). On the contrary, reweighting methods provide robust performance on target shift. Regarding covaratiate shift (Cov. shift), the improvement with reweighting methods is very limited although reweighting is specifically designed for this kind of shift. We believe that using a complex base estimator (here an SVM with an RBF kernel) enables us to train an estimator that works well on both source and target, reducing the impact of importance weighting as previously highlighted in (Byrd & Lipton, 2019) for deep neural networks. The results reported in Table 17 of Appendix E reveal that reweighting methods significantly outperform Train Src when using a linear base classifier.
Real data with unknown shift. The performance of reweighting methods is often close to Train Src baseline on real datasets. This result can be be due to the violations of the same-support assumption, which is crucial for reweighting to work effectively (Segovia-Martín et al., 2023) which is likely true for the three CV datasets. In this case, hyperparameter tuning frequently select configurations leading to near-uniform weighting, which explain the close performance to Train Src.
The performance of mapping methods is dataset-dependent, potentially due to the number of classes and presence of target shift. Mapping methods excel on MNIST/USPS and 20NewsGroup which respectively contain 10 and 2 classes, but failing on Office31 and OfficeHome with 31 and 60 classes. Additionally, while mapping performs well on the NLP dataset 20NewsGroup, it results in negative transfer on Amazon Reviews which has target shifts.
It is notable that simple transformations are the best in average across all modalities. Indeed most methods that significantly outperform Train Src in ranking average across all modalities are LinOT, CORAL, JPCA, and SA, which all rely on linear transformations such as scaling, linear projection or rotations. These methods are robust across datasets and modalities, offering effective alignment with minimal risk of negative DA.
Take-away for DA users. Reweighting methods are best suited for scenarios where the same-support assumption holds and perform particularly well when paired with regularized hypotheses like linear models. Even when assumptions are not fully satisfied, reweighting tends to be robust to negative transfer. Mapping methods are highly effective under moderate numbers of classes and in the absence of target shift but carry a significant risk of negative transfer if target shift is present. When the type of distribution shift is uncertain, simpler transformation-based methods like LinOT, CoRAL, JPCA, and SA provide modest performance improvements while minimizing risks of negative DA, making them reliable and safe default options.
@@
Selected scorer per DA method. We observe that the best scorer differs across methods, Circular Validation has been selected 10 out of 20 times as the best scorer, followed by Importance Weighting 4 out of 20 times. Table 20 in the supplementary material provides the non-realistic accuracy results with the supervised scorer. It is worth noting that the supervised scorer generally outperforms the unsupervised ones, and several methods significantly outperform Train Src in each dataset. It is crucial to choose the model realistically to avoid producing overly optimistic results, as many data analysis papers have done (see Table 4).